OpenStudy (anonymous):

How to solve this non-linear differential equation: y''-x*y'''+(y''')^3=0 @Mathematics @Calculus1

OpenStudy (anonymous):

\[y''-xy'''+(y''')^3=0\]

OpenStudy (turingtest):

That is NOT calculus 1. Though I guess there is no differential equations section...

OpenStudy (jamesj):

If the equation is with (y'')^3, then the equation is xu' - u = u^3, where u = y'' This is a Bernoulli equation and very doable. But if the equation is as written, ...

OpenStudy (anonymous):

it is as written.

OpenStudy (jamesj):

Writing as before u = y'', use an Ansatz of u = Ax + B u' = A Then as u - xu' + (u')^3 = 0, Ax + B - Ax + A^3 = 0 => B = -A^3 i.e., u = Ax - A^3 Now integrate twice to find y